1. Patients on ventilators often receive sedation to aid healing and comfort care. A doctor ordered 2.0 mcg of propofol per kilogram of body weight per minute. The pharmacy sent a 20-mcg solution. To prepare the IV, calculate how many mcg to administer per hour for a 44-lb patient.

2. A physician ordered 0.50 g of cefazolin by IV. Cefazolin exists as a 1.0 g dry package, which states to add 2.5 mL of sterile water to reconstitute to a final liquid volume of 3.0 mL. Calculate how many mL the patient should receive.
Answer #1

1. Propofol per kg - 2.0 mcg

44 lbs = 20 kgs

20 \times 2.0 = 40 mcg per minute

40 \times 60 = 2400 mcg per hour

For a patient with 44lbs, the propofol needed is 2400 mcg per hour or 2.4 mg per hour.

2. Required medication = required dose \times available volume \div available dose

                                      = 0.50 \times 3 \div 1

                                      = 1.5

To administer 0.50 gms of cefazolin by IV, we should 1.5 ml of reconstituted medication.